# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v3

@@ A.1 Proof of approximately well calibration under optimal policy
Let $R$ be the total number of answers the model can generate. Here $R$ could possibly equal infinity. Let $\hat{p}_{i}^{*}$ denote the model probability or confidence assigned to answer $i$ under optimal policy maximizing pass@N coverage in Equation 2. Let $A_{i}$ be the event that the $i$th ranked answer is not chosen in the $N$ passes and $B_{i}$ be the event that $i$th ranked answer is chosen in the $N$ passes. Let $1_{B_{i}}$ be the indicator function of $B_{i}$. Then, given an event $\omega$, the probability of getting the correct answer is $\sum_{i=1}^{R}p_{i}1_{B_{i}}(\omega)$. Hence, the expected probability of getting the correct answer is
$\mathbb{E}\left(\sum_{i=1}^{R}p_{i}1_{B_{i}}\right)=\sum_{i=1}^{R}p_{i}\mathbb{P}(B_{i})=\sum_{i=1}^{R}p_{i}(1-\mathbb{P}(A_{i}))=1-\sum_{i=1}^{R}p_{i}\mathbb{P}(A_{i})=1-\sum_{i=1}^{R}p_{i}(1-\hat{p}_{i})^{N}$ (8)
Optimal strategy maximizes Equation 8. Let $\pi_{1}(\{\hat{p}_{1}^{*},...,\hat{p}_{m}^{*}\})\equiv\hat{p}_{1}^{*}$ be the projection onto the first entry, then
$\hat{p}_{1}^{*}=\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{R}}{\text{argmin}}\{\sum_{i=1}^{R}p_{i}(1-\hat{p}_{i})^{N}\}\right).$ (9)
The proof is then motivated by the following intuition: given the constraint $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$, $\hat{p}_{1}^{*}$ attains its largest possible value when the $p_{j}\;(j\geq 2)$ are as small as possible. It is worth noting that first
$\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{R}}{\text{argmin}}\{\sum_{i=1}^{R}p_{i}(1-\hat{p}_{i})^{N}\}\right)\leq\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right).$ (10)
Second, given $p_{1}$ and $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$, $\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right)$ is bounded from above by
$\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+\sum_{i=2}^{k}\frac{1-p_{1}-\epsilon}{k-1}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right).$ (11)
Third, fix $\hat{p}_{1}$, Jensen’s inequality tells us that
$p_{1}(1-\hat{p}_{1})^{N}+\sum_{i=2}^{k}\frac{1-p_{1}-\epsilon}{k-1}(1-\hat{p}_{i})^{N}\geq p_{1}(1-\hat{p}_{1})^{N}+(1-p_{1}-\epsilon)(1-\frac{1-\hat{p}_{1}}{k-1})^{N}.$ (12)
Combining all three inequalities above we have
$\hat{p}_{1}^{*}=\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{R}p_{i}(1-\hat{p}_{i})^{N}\}\right)\leq\underset{\hat{p}_{1}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+(1-p_{1}-\epsilon)(1-\frac{1-\hat{p}_{1}}{k-1})^{N}\}.$ (13)
The right hand side of Equation 13 can be computed by taking derivatives, and we arrive at
$\hat{p}_{1}^{*}&\leq\underset{\hat{p}_{1}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+(1-p_{1}-\epsilon)(1-\frac{1-\hat{p}_{1}}{k-1})^{N}\}\\ &=1-\frac{k-1}{(k-1)^{\frac{N}{N-1}}p_{1}^{\frac{1}{N-1}}\left(1-p_{1}-\epsilon\right){}^{\frac{1}{1-N}}+1}\end{aligned}.$ (14)
@@
Computing the right hand side of Equation 17 by taking derivatives we have
@@
Combining Equation 19 and Equation 18 we arrive at
@@ Appendix B Experimental details
Our codebase uses PyTorch [45], Accelerate [46], and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm [47] for efficient inference. Code will be made available at https://github.com/allanraventos/refine.
MATH. For experiments with MATH dataset [41] (license: MIT), we fine-tune the Llama-3-8B-base [40] on the MATH [41] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. [22] and use $12,000$ problems for training and the remaining $500$ for testing. In Sections 4 and 5.2, Figure 1, Figure 2 (a) and (b), we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section 5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
@@
DCOa objective. The DCOa introduced in Section 5.4 is an approximation for DCO. To construct a batch of size $B$ with DCOa, we process batches of samples sequentially; for each batch, we run online inference on each of the samples and discard all samples with probability of success rate larger than $p^{\mathrm{thresh}}$. We choose to discard samples which have a DCO confidence reguarizer $F$ lower than 0.01, corresponding to $p^{\mathrm{thresh}}=0.1$ for $N^{\prime}=64$. This process continues until we have enough training data for a single batch. In Figure 4, we plot the number of discarded samples as a function of training step for our DCOa experiments (same ones as in Table 3).
@@
Implementing online inference. Throughout our experiments and analysis, we extensively use the open-source vllm package [47] for efficient inference. Integrating inference into the training loop to enable training under the DCOa objective, as in Section 5.4, poses a challenging implementation problem. We solve this problem by placing vllm worker processes, which together perform inference on all GPUs, in a separate process group and use Ray (https://github.com/ray-project/ray) to isolate them from the training loop. This enables running concurrent training and inference on the same set of GPUs. We believe this inference in the loop setup will be useful to the community, as it enables straightforward implementation of online data filtering approaches for LLM training.
@@ C.1 Easy data drives overconfidence
Figure 5: Easy data drives overconfidence and degrades performance when scaling test-time compute. For the experiments in Figure 1, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ at two points during the first training epoch: 11% (early) and 86% (late). At each stage, the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$ is evaluated on batches of previously unseen training data of each difficulty level defined in the MATH dataset. Early in training, data from all difficulty levels contribute to decreasing the test loss (left). However, later in training, easier examples (difficulty level 2) provide no further benefit, while the easiest examples (difficulty level 1) actively degrade test performance (right). The plotted $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is an average over batches of unseen data, and we use $N=256$, corresponding to pass@256.
We have shown in Section 4.2 that overconfidence limits performance gains from scaling test-time compute. Here we investigate whether all training data contribute equally to the drop in pass@N test performance. To quantify this effect, we define the test loss as the negative log-probability of coverage, $\mathcal{L}_{\text{test}}^{N}=-\sum_{(x,y)\in\mathcal{D}^{\mathrm{test}}}\mathcal{\log}\ \mathcal{C}_{(x,y)}^{N}$, where $\mathcal{C}_{(x,y)}^{N}$ denotes the coverage for a single test example $(x,y)$. To measure how a training batch $\mathcal{B}$ influences the final test loss, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ for the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$, where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example. A negative value, $\nabla_{g}\mathcal{L}_{\text{test}}^{N}<0$, indicates that a gradient step in the direction of $g$ decreases the test loss.
To analyze the impact of data difficulty, we group training data by difficulty level (as specified in the MATH dataset) and use the directional derivative to examine how a batch of previously unseen data from a given difficulty level would affect $\mathcal{L}_{\text{test}}^{N}$ at different stages of training. Early in training, we observe that $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is negative across all difficulty levels, indicating that data from all difficulty levels contribute positively to reducing the test loss (Figure 5, left). However, near the end of the first epoch, training on easier examples (difficulty levels 1 and 2) no longer improves performance, and the easiest examples (level 1) actively degrade pass@N test performance (Figure 5, right). This shows how continued training on easy data can harm test performance when scaling test-time compute.
@@
Figure 6: Pareto-optimality at a given test strategy pass@N for some $N$ is obtained by DCO training for some $N^{\prime}$ close to $N$. Same as Figure 1 but with $N^{\prime}=256$ highlighted. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. Note that there is no choice of $N^{\prime}$ that is optimal across all $N$ (different colors are higher at different $N$). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Pareto-optimality at a given test strategy pass@N for some $N$ on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.3 Addition results with Llama-3-70B-base

### v4

@@ A.2.1 Case $k$
In this case, we have
$\hat{p}_{k}^{*}=1-\frac{(k-1)p_{k}^{-\alpha}}{\sum_{s=1}^{k}p_{s}^{-\alpha}}>0,\;\hat{p}_{1}^{*}=1-\frac{(k-1)p_{1}^{-\alpha}}{\sum_{s=1}^{k}p_{s}^{-\alpha}}.$ (10)
where $\alpha=\frac{1}{N-1}$. Suppose, for contradiction, that $p_{1}>p_{k-1}$. We construct a perturbed configuration $\{q_{i}\}_{i=1}^{k}$ such that the corresponding optimizer $\{\hat{q}_{i}^{*}\}_{i=1}^{k}$ satisfies $\hat{q}_{1}^{*}>\hat{p}_{1}^{*}$. Choose a small $\delta>0$ and define
$q_{1}=p_{1},\qquad\dots\qquad,q_{k-2}=p_{k-2},\qquad q_{k-1}=p_{k-1}+\delta,\qquad q_{k}=p_{k}-\delta.$
For $\delta>0$ sufficiently small, we still have $\hat{q}_{k}^{*}>0$ (by continuity), and moreover $\hat{q}_{1}^{*}>\hat{p}_{1}^{*}$, contradicting the optimality of the original configuration $\{p_{i}\}_{i=1}^{k}$. Hence, $p_{1}=p_{2}=\dots=p_{k-1}$. ∎
By Lemma A.2, we have
$(k-1)p_{1}+p_{k}=1-\epsilon,\;(k-1)p_{1}^{-\alpha}>(k-2)p_{k}^{-\alpha},\;p_{1}\geq p_{k}>0.$ (11)
This implies the following range for $p_{1}$ in Case $k$:
$\frac{1-\epsilon}{k}\leq p_{1}<\frac{1-\epsilon}{k-1+\left(\frac{k-2}{k-1}\right)^{N-1}}.$ (12)
Under condition Equation˜12, we have
$u=1-\frac{(k-1)p_{1}^{-\alpha}}{(k-1)p_{1}^{-\alpha}+(1+p_{1}-\epsilon-kp_{1})^{-\alpha}}.$ (13)
Moreover, the right hand side of Equation˜13 is an increasing function of $p_{1}$.
A.2.2 Case $j$
Lemma A.3.
If $\hat{p}_{j}^{*}>0$ but $\hat{p}_{j+1}^{*}=0$ for some $2\leq j\leq k-1$, then $p_{1}=...=p_{j-1}$ and $p_{j+1}=...=p_{k}>0$.
Proof.
We proceed in three steps.
Step 1: we prove $p_{k}>0$. Suppose, by contradiction, $p_{k}=0$. Let $s>2$ be the smallest index such that $p_{s}=\dots=p_{k}=0$. If $s>j$, change $p_{j}$ to $p_{j}-\delta$, $p_{j+1}$ to $\delta$ for some small $\delta>0$ If $s\leq j$, change $p_{s-1}$ to $p_{s-1}-\delta$, $p_{j}$ to $\delta$ for some small $\delta>0$. In either case, for $\delta$ sufficiently small, the new system remains in Case $j$ but has strictly larger $\hat{p}_{1}^{*}$, a contradiction.
Step 2: we prove $p_{j+1}=\dots=p_{k}$. Being in Case $j$ implies
$c_{j}>0,$ (14)
$c_{r}\leq 0,\;\forall j+1\leq r\leq k,$
$\text{where}\;c_{r}\equiv 1-\frac{(r-1)p_{r}^{-\alpha}}{\sum_{t=1}^{r}p_{t}^{-\alpha}}.$
If $c_{j+1}<0$, change $p_{j}$ to $p_{j}-\delta$, $p_{j+1}$ to $p_{j+1}+\delta$ for some small $\delta>0$. The new system is still in Case $j$ but with strictly larger $\hat{p}_{1}^{*}$, which leads to contradiction. Hence $c_{j+1}=0$.
If $c_{j+2}<0$, then $p_{j+2}<p_{j+1}$ and we can change $p_{j+1}$ to $p_{j+1}-\delta$, $p_{j+2}$ to $p_{j+2}+\delta$ for some small $\delta>0$. Under this change, $c_{j+1}$ decreases, $c_{j+2}$ increases, $c_{s},s\geq j+3$ decreases, so the configuration remains in Case $j$ but with equal $\hat{p}_{1}^{*}$ for $\delta>0$ sufficiently enough. But then $c_{j+1}<0$, contradicting $c_{j+1}=0$. Iterating this argument yields $c_{j+1}=\dots=c_{k}=0$, which implies $p_{j+1}=\dots=p_{k}$.
Step 3: We prove $p_{1}=\dots=p_{j-1}$. This follows from the same argument as in the proof of Lemma˜A.2. ∎
As a direct consequence, we obtain the following corollary.
Corollary A.4.
Let $p_{j}=q$, $p_{j+1}=...=p_{k}=p$. The upper bound $u$ is given by
$u=1-\frac{(j-1)p_{1}^{{}^{-\frac{1}{N-1}}}}{(j-1)p_{1}^{-\frac{1}{N-1}}+q^{-\frac{1}{N-1}}}=\frac{q^{-\frac{1}{N-1}}}{(j-1)p_{1}^{-\frac{1}{N-1}}+q^{-\frac{1}{N-1}}}.$ (15)
Moreover,
$\frac{1}{j}\leq u<\frac{1}{j-1}.$ (16)
and $\{q,p,j\}$ satisfies
$p_{1}\geq q\geq p>0,$ (17)
$(j-1)p_{1}+q+(k-j)p=1-\epsilon,$
$(j-1)(p^{-\alpha}-p_{1}^{-\alpha})=q^{-\alpha},\;\alpha=\frac{1}{N-1},$
$q>\left(\frac{j-2}{j-1}\right)^{N-1}p_{1}.$
Proof.
This is an immediate consequence of Lemma˜A.3 and its proof (using $c_{j+1}=0$, $c_{j}>0$). ∎
Once $j$ is determined, we can compute $q$ and hence $u$. Equation˜17 determines $j$ in the following sense.
Lemma A.5.
If $\{q_{0},p_{0},j_{0}\}$ satisfies Equation˜17, then $j_{0}\geq j$.
Proof.
Let $u_{0}$ be the corresponding maximal value associated with $\{q_{0},p_{0},j_{0}\}$. Then $u_{0}\leq u$. Combining Equation˜17 and Equation˜15 we have
$\frac{1}{j_{0}}\leq u_{0}<\frac{1}{j_{0}-1}.$ (18)
If $j<j_{0}$, then $u_{0}>u$, which is a contradiction. ∎
Lemma A.6.
For any fixed $\epsilon$, $p_{1}$ and $k$, $j=j(N)$ is a non-decreasing function of $N$. Since $1/j\geq u>1/(j-1)$, $u=u(p_{1})$ is a non-decreasing function of $N$.
Proof.
Suppose, for contradiction, that $j(N+1)<j(N)$. Then
$\left(\frac{j(N+1)-2}{j(N+1)-1}\right)^{N}<\left(\frac{j(N)-2}{j(N)-1}\right)^{N-1}.$ (19)
Let $\{j(N+1),q(N+1),p(N+1)\}$ be the corresponding configuration at $N+1$. Using the inequality above, this configuration can be used to construct a larger value of $u$ at $N$ with index $j=j(N+1)<j(N)$, contradicting the minimality of $j$ in Lemma A.5. ∎
A.2.3 Case 1
This case is impossible.
A.2.4 Summary of the Proof of Lemma 4.2
  • If
    $\frac{1-\epsilon}{k}\leq p_{1}<\frac{1-\epsilon}{k-1+\left(\frac{k-2}{k-1}\right)^{N-1}},$ (20)
    then Case $k$ applies. In this regime, the upper bound $u$ is an increasing function of $p_{1}$ and is given by
    $u=1-\frac{(k-1)p_{1}^{-\frac{1}{N-1}}}{(k-1)p_{1}^{-\frac{1}{N-1}}+(1+p_{1}-\epsilon-kp_{1})^{-\frac{1}{N-1}}}.$ (21)
  • Otherwise, Case $j$ applies, and the upper bound is
    $u=1-\frac{(j-1)p_{1}^{{}^{-\frac{1}{N-1}}}}{(j-1)p_{1}^{-\frac{1}{N-1}}+q^{-\frac{1}{N-1}}}.$ (22)
    For any fixed $\epsilon$, $p_{1}$ and $k$, let $j$ be the smallest positive integer such that the following equations (viewing $\{q,p\}$ as variables) admits a solution:
    $p_{1}\geq q\geq p>0,$ (23)
    $(j-1)p_{1}+q+(k-j)p=1-\epsilon,$
    $(j-1)(p^{-\alpha}-p_{1}^{-\alpha})=q^{-\alpha},\;\alpha=\frac{1}{N-1}.$
    $q>\left(\frac{j-2}{j-1}\right)^{N-1}p_{1}.$
    The index $j$ is non-decreasing with respect to $N$, and $u$ is further bounded by
    $\frac{1}{j}\leq u<\frac{1}{j-1}.$ (24)
@@ A.2.3 Case 1
Computing the right hand side of Equation˜27 by taking derivatives we have
@@
Combining Equation˜29 and Equation˜28 we arrive at
@@ Appendix B Experimental details
Our codebase uses PyTorch [47], Accelerate [48], and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm [49] for efficient inference. Code will be made available here (https://github.com/allanraventos/refine).
MATH. For experiments with MATH dataset [43] (license: MIT), we fine-tune the Llama-3-8B-base [42] on the MATH [43] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. [24] and use $12,000$ problems for training and the remaining $500$ for testing. In Sections˜4 and 5.2, Figure˜1, Figure˜2 (a) and (b), we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section˜5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
@@
DCOa objective. The DCOa introduced in Section˜5.4 is an approximation for DCO. To construct a batch of size $B$ with DCOa, we process batches of samples sequentially; for each batch, we run online inference on each of the samples and discard all samples with probability of success rate larger than $p^{\mathrm{thresh}}$. We choose to discard samples which have a DCO confidence reguarizer $F$ lower than 0.01, corresponding to $p^{\mathrm{thresh}}=0.1$ for $N^{\prime}=64$. This process continues until we have enough training data for a single batch. In Figure˜4, we plot the number of discarded samples as a function of training step for our DCOa experiments (same ones as in Table˜3).
@@
Implementing online inference. Throughout our experiments and analysis, we extensively use the open-source vllm package [49] for efficient inference. Integrating inference into the training loop to enable training under the DCOa objective, as in Section˜5.4, poses a challenging implementation problem. We solve this problem by placing vllm worker processes, which together perform inference on all GPUs, in a separate process group and use Ray (https://github.com/ray-project/ray) to isolate them from the training loop. This enables running concurrent training and inference on the same set of GPUs. We believe this inference in the loop setup will be useful to the community, as it enables straightforward implementation of online data filtering approaches for LLM training.
@@ C.1 Easy data drives overconfidence
Figure 5: Easy data drives overconfidence and degrades performance when scaling test-time compute. For the experiments in Figure˜1, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ at two points during the first training epoch: 11% (early) and 86% (late). At each stage, the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$ is evaluated on batches of previously unseen training data of each difficulty level defined in the MATH dataset. Early in training, data from all difficulty levels contribute to decreasing the test loss (left). However, later in training, easier examples (difficulty level 2) provide no further benefit, while the easiest examples (difficulty level 1) actively degrade test performance (right). The plotted $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is an average over batches of unseen data, and we use $N=256$, corresponding to pass@256.
We have shown in Section˜4.2 that overconfidence limits performance gains from scaling test-time compute. Here we investigate whether all training data contribute equally to the drop in pass@N test performance. To quantify this effect, we define the test loss as the negative log-probability of coverage, $\mathcal{L}_{\text{test}}^{N}=-\sum_{(x,y)\in\mathcal{D}^{\mathrm{test}}}\mathcal{\log}\ \mathcal{C}_{(x,y)}^{N}$, where $\mathcal{C}_{(x,y)}^{N}$ denotes the coverage for a single test example $(x,y)$. To measure how a training batch $\mathcal{B}$ influences the final test loss, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ for the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$, where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example. A negative value, $\nabla_{g}\mathcal{L}_{\text{test}}^{N}<0$, indicates that a gradient step in the direction of $g$ decreases the test loss.
To analyze the impact of data difficulty, we group training data by difficulty level (as specified in the MATH dataset) and use the directional derivative to examine how a batch of previously unseen data from a given difficulty level would affect $\mathcal{L}_{\text{test}}^{N}$ at different stages of training. Early in training, we observe that $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ is negative across all difficulty levels, indicating that data from all difficulty levels contribute positively to reducing the test loss (Figure˜5, left). However, near the end of the first epoch, training on easier examples (difficulty levels 1 and 2) no longer improves performance, and the easiest examples (level 1) actively degrade pass@N test performance (Figure˜5, right). This shows how continued training on easy data can harm test performance when scaling test-time compute.
@@
Figure 6: Pareto-optimality at a given test strategy pass@N for some NN is obtained by DCO training for some $N^{\prime}$ close to $N$. Same as Figure˜1 but with $N^{\prime}=256$ highlighted. We fine-tune Llama-3-8B base models on the MATH dataset to produce direct answers. We fine-tune for 4 epochs one model using CE loss and several models under the $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ objective, for choices of $N^{\prime}$ indicated by color. Note that there is no choice of $N^{\prime}$ that is optimal across all NN (different colors are higher at different NN). The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. Fine-tuning with the DCO loss ($N^{\prime}=256$) achieves Pareto-optimality for $180\leq N\leq 385$ at test-time, further confirming that Pareto-optimality at a given test strategy pass@N for some NN on the x-axis is obtained by DCO training for some $N^{\prime}$ close to $N$.
C.3 Additional results with Llama-3-70B-base